Solve the following recurrence relation without using the master method! report the big O

1. T(n) = 2T(n/2) =n^2

2. T(n) = 5T(n/4) + sqrt(n)
Answer #1


T(n)= at(M) + O(nd) a=hd Ти) ә0 (и tolnd) de O Inlogg) -- a<6d a> bd 1. Tul= 21(4) & u² a=2 b=2 d = 2 6 = 4 d acod T(nl= O(n)